OpenStudy (anonymous):

If (-3, y) lies on the graph of y = (1/2)x, then y =

OpenStudy (anonymous):

given the point. X=-3, right?

OpenStudy (anonymous):

so plug x into y= (1/2)x y = (1/2)(-3) y = -1.5

OpenStudy (anonymous):

If (-3, y) lies on the graph of y = (1/2)x, then y =? The answer to this math question is plot | y = 1/2 x.

OpenStudy (tkhunny):

Don't make multiple posts. http://openstudy.com/study#/updates/5351b06be4b0b39169ce0fd6